Assume the Honda Motor Company sells the Honda Accord in the United States for $23,175. a. If the markets are efficient, how much should the Honda Accord sell for in Great Britain if the exchange rate is $1.80/£? b. How much would the price of the Honda Accord change in Great Britain if the U.S. dollar were to appreciate to an exchange rate of $1.65 next year?
Answer #1

Price of car in UK = price in USD/exchange rate

=23175/1.8=12875

New price in UK = price in USD/exchange rate

=23175/1.65=14045.4545

Change in price = new price-old price = 14045.4545-12875=1170.4545